Calculate the number Of ions in 500cm3 of 0.5mol/dm3 (NH4)2Cu(SO4)2(aq)?
1
Expert's answer
2020-11-26T05:08:33-0500

n((NH4)2Cu(SO4)2) = C*V = 0.5mol/dm3*0.5dm3 = 0.25 mol

(NH4)2Cu(SO4)2 = 2NH4^+ + Cu^2+ + SO4^2-

According to the equation

n(ions) = 4*0.25 mol = 1 mol

Number of ions - 6.02*10^23 ions